# Ultra-High-Energy Cosmic Rays from Neutrino-Emitting Tidal Disruption Events

### v2

@@ 3.1 Black hole mass function and TDE rates
The cosmological rate of TDEs can be estimated as the product of three functions: the rate of TDEs per black hole , the Super-Mass Black Hole (SMBH) mass function – defined as the number of black holes per comoving volume and per unit mass at redshift – and the occupation fraction, , which is the probability that a SMBH is located at the center of a host galaxy: . The mass function scales roughly as , and like for all and for (Shankar et al., 2009). The fraction can be approximated as a step function, with () above (below) a cutoff mass , which represents the threshold below which suppressing effects, like black hole ejection from the host galaxy come into play (see, e.g., Fialkov & Loeb (2017)). The value of is uncertain, with different estimates ranging in the interval (e.g. Stone & Metzger, 2016; Kochanek, 2016; Fialkov & Loeb, 2017). The rate of tidal disruptions per SMBH is estimated to be (Kochanek, 2016), where the dependence on the type of disrupted star is neglected in first approximation.
For simplicity, here, for each class of TDE (e.g. TDEs involving main sequence stars, or White Dwarfs, etc.), we use the total cosmological rate of TDEs (number of TDEs per comoving volume per unit time) as a function of , .
In particular, we take , following the evolution of the number density of SMBHs as calculated in Shankar et al. (2009) (see also Stone & Metzger (2016); Kochanek (2016); Lunardini & Winter (2017)). The normalization, , is highly uncertain because it depends on several effects. For main sequence stars, a major source of uncertainty is the interval of SMBH mass, , for which tidal disruption is efficient; in particular, varying the lower mass cutoff can cause an uncertainty of about an order of magnitude: (see Lunardini & Winter (2017)), whereas the effect of the upper cutoff, – the mass beyond which a star is swallowed whole and not disrupted — is smaller. For post-main-sequence stars, rates are typically lower. For example, the rate is estimated at for red giants (Syer & Ulmer, 1999), and should be orders of magnitude smaller for red supergiants, which are a lot more rare (a specific rate estimate could not be found for these).
TDEs involving a white dwarf star can occur only for Galactic black holes in the lowest mass range (“intermediate” mass BHs, ) and area significantly rarer: , see, e.g. Krolik & Piran (2011); Senno et al. (2017); Zhang et al. (2017).
@@ 3.2 From two to many: Extrapolating the model to the TDE population
A key question here is how to assign values of the fitted parameters and other relevant input parameters (e.g., IR photon temperature and luminosity, etc.) to each member of the UHECR-neutrino emitting TDE population. We are taking a model-independent approach and assuming that each of the AT2019aalc and AT2019fdr represents its own unique group. We will refer to them as the “aalc-group” and the “fdr-group”, which together represent the total population. Each group has its unique parameters and local rate (, where aalc or fdr). The local rate of the entire UHECR-neutrino emitting TDE population is given by . The analyzed range and values of these parameters are summarized in Tab. 3
@@
Considering the large number of fitting variables and the limited amount of data, we have to consider different scenarios, which are compared to each other. In particular, we examine the following scenarios:
@@
For each scenario, we consider two options for the composition of the injected CR, as discussed in Sec. 2.3. The first option follows the maximum abundance of a star from the selected statistics (“Star Mass Abundance”). The second option treats the composition of the injected CR as fitted parameters (“Fitted Composition Case”).
@@ 3.3 UHECR transport to Earth and description of UHECR data
In our analysis of UHECRs originating from TDEs, we employ the open-source software PriNCe (Heinze et al., 2019), designed to numerically solve the one-dimensional transport equations, to model the propagation of CRs in the Universe and compute the cumulative cosmogenic neutrino spectra. We propagate the CRs escaped from the source for each aalc- and fdr-group, considering varying radius, maximal rigidity, and injected composition. We consider the redshift evolution up to for the TDE populations to calculate the cumulative diffuse neutrino fluxes. For the configuration of the calculation, we use the build-in Talys (Koning et al., 2007) nuclear interaction model, and use the Extragalactic Background Light (EBL) model by Gilmore et al. (2012).
Our primary goal is to fit the UHECR spectrum (Verzi, 2019) and composition (Yushkov, 2019) data above , observed by the Pierre Auger Observatory (Auger). In addition, we ensure that the diffuse neutrino flux from the TDE population does not exceed the High-Energy Starting Events (HESE) data from IceCube. For the composition fitting, we attempt to reproduce the observed mean depth of shower maximum density () and its standard deviation (), using the EPOS-LHC air-shower model (Pierog et al., 2015)333We also tested the Sibyll 2.3d air-shower model (Riehn et al., 2020) and found no significant difference. The results from Sibyll 2.3d air-show model are available at plotkopavlo.com/uhecrsources/tde.. To account for systematic, we include the energy and uncertainties. However, we exclude the uncertainties since they have a minor influence on our analysis.
To evaluate the goodness of the fitting, we adopt the method from Plotko et al. (2023) and calculate the value for the corresponding model of the TDE population:
@@
Further, to compare different models of the TDE population, we compute the Akaike Information Criterion corrected (, Akaike, 1974; Burnham & Anderson, 2004; Buchner, J. et al., 2014; Rosales et al., 2020) through the given equation:
where is the number of data points, and is the number of parameters of the corresponding model . The model with the smallest value explains the data best among tested models. It is important to highlight that should not be used as a null-hypothesis test (Anderson et al., 2000; Mundry, 2011). In our study, we compare the score of each model against of the model with the smallest value (), calculating the difference () as:
To make it easier to observe the differences between models, we compute in the units of standard deviations following Plotko et al. (2023), and use it as an identifier for model selections.
@@ 4.1 Evaluation of TDE population models
Scenario | Composition | Group’s local rate | Fit quality metrics | ||||
aalc-group | fdr-group | aalc-group | fdr-group | / d.o.f. | |||
Mono | |||||||
aalc | MS | 10 | 654/29 | 603 | |||
CO-WD | 59 | 84/29 | 33 | 5.3 | |||
ONeMg-WD | 75 | 192/29 | 140 | ||||
Fitted | 122 | 57/27 | 12 | 2.8 | |||
fdr | Fitted | 9 | 261/27 | 216 | |||
Democratic | |||||||
MS | 69 | 37650/29 | 37599 | ||||
CO-WD | 75 | 2420/29 | 2369 | ||||
ONeMg-WD | 69 | 447/29 | 396 | ||||
Fitted | 69 | 248/27 | 203 | ||||
Diverse | |||||||
ONeMg-WD | WR | 47 | 5 | 57/24 | 23 | 4.3 | |
Fitted | Fitted | 194 | 13 | 38/25 | Best-fit case |
@@ 4.2 Describing the UHECR spectra and composition with TDE population
In this section, we present results for the best-fit case of the UHECR-neutrino emitting TDE population, which is the Diverse Scenario in the Fitted Composition Case. In this scenario, the aalc- and fdr-groups have different parameters, such as the radius of the production region, the maximum rigidity, and the injected CR composition. Fig. 6 illustrates the results for UHECR spectrum and composition and diffuse neutrino flux in comparison to data for the best-fit model. The obtained fitting parameters for the Diverse Scenario are detailed in Tab. 5.
@@
Our analysis shows that the fitted UHECR spectrum closely aligns with the spectrum observed by the Auger Observatory, with the fitted value of the systematic energy shift of about -9.4%. The aalc-group primarily dominates the UHECR spectrum. Although the fdr-group’s overall contribution is subdominant in spectral fitting, it plays a crucial role in composition fitting, particularly at the highest energies. The model adjustment results in a heavier composition by shifting by approximately 174%. Notably, this adjustment is consistent with recent findings from Auger, e.g., a potential systematic bias toward a heavier composition (Abdul Halim et al., 2024b).
@@
We note that while the UHECR are fitted, the predicted diffuse neutrino flux is a consequence from that fit; it is quite large, but does not exceed IceCube’s HESE data. The contribution becomes significant at energies of GeV and higher, peaking at GeV. Unlike the UHECR spectrum, both the aalc- and fdr- groups contribute at a comparable level to the diffuse neutrino flux (see the upper right panel of Fig. 6). However, we also observe that the aalc-group produces a more peaked spectrum with a lower maximum energy in contrast to the fdr-group. As a result, the two groups predict different slopes of the neutrino spectrum in the EeV energy range. As can be seen on the upper right panel of Fig. 6), both groups are likely detectable by forthcoming neutrino detectors like IceCube-Gen2 and GRAND200k. Regrettably, the predicted diffuse flux is too low for RNO-G to detect. As far as the peak energy of the neutrino flux is concerned, it is driven by the primary energies and considerably higher than the originally claimed observations in the 80–220 TeV range; a possible reason for this difference is that the IceCube detector response function is typically based on power law spectra, whereas peaky spectrum lead to systematically higher reconstructed neutrino energies, see Rodrigues et al. (2024) for AGN blazars.
@@
Regarding the composition of injected UHECR, the aalc-group emits CRs in both intermediate-mass (e.g., sodium or silicon) and the light (e.g., hydrogen and helium) ranges, whereas the fdr-group mostly contributes the heavy nuclei in the best-fit scenario. This distribution is depicted in Fig. 7 and detailed in Tab. 5. The composition fitting for the aalc-group is similar to ONeMg-WD. Still, it requires additional light elements. The absence of intermediate elements could be attributed to the loss of intermediate elements before they are transferred to the acceleration zone. Furthermore, more iron is needed for the fdr-group, to make the predicted composition heavier to fit the composition data better. However, within the 1 AICc region, the injected composition could be pure silicon, which is similar to the ONeMg-WD. One way to achieve an iron-abundant CR injection for both groups is by disrupting a MS star with an enhancement factor for the acceleration of heavier elements, which will be discussed in the Sec. 5.
@@
In order to match the observed flux of UHECR, a local rate of 200 is required for the aalc-group. This value is typical for TDEs that disrupt MS stars rather than white dwarfs. For the fdr-group, the local rate should be considerably lower, around 10 .
Further, Fig. 8 explores the parameter spaces for the aalc- and fdr-groups, and Fig. 9 presents the escaped spectra of CRs and neutrinos for the best-fit parameters. The aalc-group parameters are notably similar to those previously modeled for AT2019aalc (Winter & Lunardini, 2023), with a maximum rigidity value of GV that is expected for UHECR sources (Heinze et al., 2019; Aab et al., 2017; Plotko et al., 2023). The composition 3 region requires values between and GV to fit the data. Higher values are excluded since, in this case, the escape CRs from TDE are not dominated by a single element group but also have too many secondary light elements, contradicting the data.
On the other hand, the 3 allowed region, based on the spectral fitting, for the aalc-group (marked as the yellow region in the left panel of Fig. 8) is broad and does not significantly impact the fit. The spectrum is less sensitive to composition and does not provide a clear reason for this shape of the 3 spectrum allowed region. The required value of the TDE source radius of the aalc-group is cm, falling within the purple region allowed by the neutrino constraints for AT019aalc.
Conversely, the fdr-group requires substantially different parameters from previously used in Winter & Lunardini (2023). In order to fit the composition data, the fdr-group should have a higher maximum rigidity, with the best value of GV – which was, however an ad hoc choice in the original paper to surpass the pion production threshold, and should rather be understood as a lower limit there. This suggests that TDEs from the fdr-group could be a powerful UHECR source capable of accelerating iron to energies of GeV. The preference for a high maximum rigidity arises because iron nuclei are required to contribute at energies around GeV, while their presence at lower energies should be negligible. Spectral constraints are not applied to the fdr-group since its overall contribution to the UHECR spectrum is too low to significantly impact the fit.
@@ 5 Enhancement of heavier nuclei in the acceleration process
In order to further investigate the possibility of obtaining a heavier CR composition, we introduce an enhancement factor for the acceleration of heavier elements. Here, the idea is to transfer an over-proportionally large fraction of energy into heavier non-thermal nuclei compared to what would be naively expected from the mass composition of the progenitor star. Realistically, nucleus selection effects may exist and cause the composition of the accelerated nuclei to differ from that of the disrupted star. One such effect might be partial disruption, which is supported by recent observations (Somalwar et al., 2023). In this case, the core of the star would remain intact while the matter in the outer layers is disrupted and subsequently accreted onto the central black hole. Ultimately, the accreted material could be accelerated, resulting in the composition of the observed UHECRs. Another possibility is that there might be enhancement mechanisms in the acceleration process that would favor heavier nuclei, thus skewing the UHECR toward heavier composition compared to the original star.
We define an enhancement factor based on Caprioli et al. (2017) and Hanusch et al. (2019), as Caprioli et al. (2017) showed that nonrelativistic collisionless shocks can preferentially accelerate ions based on their mass/charge ratio ( ratios). As a result, these shocks develop a non-thermal tail whose normalization is enhanced as . Hanusch et al. (2019) further supports the acceleration of heavier ions through hybrid simulations, which reveal that the injection efficiency of ions into the diffusive shock acceleration process depends strongly on the shock Mach number. Their simulations showed that the injection efficiency for ions with high ratios is enhanced. To generalize the idea, we assume that injection efficiency is proportional to . Additionally, we assume that thermal particles lose an electron () before being transferred to an acceleration zone to maximize the effect. The net effect is that the abundance of the element in the accelerated stream is scaled by a power of the mass number, leading to new, rescaled mass fractions, :
@@
In our analysis, we use the Diverse Scenario and focus on the fixed composition of the MS star with various values of for injected CRs. In this scenario, we assume that the aalc- and fdr-groups have different values. For each combination of and , we conduct a fit to determine the minimum value while minimizing over all other fitted parameters. We show as a function of the enhancement factor in Fig. 11. For the aalc-group, the fit requires the value of to be between 1.5 and 2. For the fdr-group, there is a slight preference for values of for the fdr-group, but not statistically significant due to the small contribution of the fdr-group to the UHECR spectrum.
Overall, the Enhancement Factor Case is excluded by more than 8 compared to the Fitted Composition Case due to the nature of the enhancement factor. The model predicts a light composition for values of the enhancement factor near zero, which contradicts the UHECR composition observations. As the value of increases, the model becomes too efficient at accelerating heavy elements. However, to explain UHECR data, an intermediate-mass group is needed. Different assumptions could help address this issue. For example, there may be two groups for TDEs: one that disrupts white dwarfs with intermediate mass elements and another that disrupts MS stars with an enhancement factor for acceleration of heavier elements. This approach would provide both lighter and heavier elements, potentially improving the fit of the composition data. However, such a model would introduce too many free parameters and could be statistically excluded in favor of the simpler Fitted Composition Case.
Fig. 12 compares the best-fit case to the Auger spectrum and composition, along with the predicted diffuse neutrino flux, for the Enhancement Factor Case. The corresponding parameters of the best-fit case are presented in Tab. 6. While the model’s predictions can explain the UHECR spectrum, the main contribution comes from the iron group, which contradicts the current UHECR composition observation. Additionally, the diffuse neutrino flux from the aalc-group exceeds the HESE limits with a high local rate of 1116 . These two factors result in a poor quality of fit. One potential solution to improve the fit could be using a more complex enhancement factor, such as having iron elements lose more electrons than the silicon group, resulting in a more silicon-dominant composition.

### v3

@@ 3.1 Black hole mass function and TDE rates
For each class of TDE (i.e., TDEs involving MS stars, WD, etc.), we use the total cosmological rate of TDEs (number of TDEs per comoving volume per unit time) as a function of , of the form . The redshift dependence reflects the (negative) cosmological evolution of the SMBH and IMBH mass function, as described in, e.g. F. Shankar et al. (2009) (see also N. C. Stone & B. D. Metzger (2016); C. S. Kochanek (2016); C. Lunardini & W. Winter (2017)). The normalization, , depends on the details of the galactic population of stars and of SMBH/IMBH, and in particular on the disruption rate per black hole, , and on the black hole occupation fraction, (the probability that a SMBH or IMBH is located at the center of a host galaxy).
For main sequence star disruptions, a major source of uncertainty is the interval of SMBH mass, , for which tidal disruption is efficient: varying the lower mass cutoff can cause an uncertainty of about an order of magnitude: (see C. Lunardini & W. Winter (2017)), whereas the effect of the upper cutoff, – the mass beyond which a star is swallowed whole and not disrupted — is smaller.
Disruption of white dwarfs must be significantly rarer, due to several factors. One of these is the smaller tidal disruption radius: in Ref. J. H. Krolik & T. Piran (2011) it is reasoned that - if a similar number density of WD and MS stars is assumed - is suppressed by the ratio of the tidal disruption radii: . Further suppression may come from the flattening of the BH mass function at (see discussion in M. MacLeod et al. (2016) and recent results in H. Cho & J.-H. Woo (2024)), and from the possibly low (and uncertain) occupation fraction for IMBH (H. Cho & J.-H. Woo, 2024). Considering the large uncertainties, rates in the range , are realistic, see, e.g. J. H. Krolik & T. Piran (2011); R. V. Shcherbakov et al. (2013); B. T. Zhang et al. (2017). Extreme values, up to are found using recent numerical simulations (A. Tanikawa et al., 2022).
For post-main-sequence stars, rates are typically lower than for MS stars. For example, the rate is estimated at for red giants (D. Syer & A. Ulmer, 1999), and should be orders of magnitude smaller for red supergiants, which are a lot more rare (a specific rate estimate could not be found for these).
@@ 3.2 From two to many: Extrapolating the model to the TDE population
A key question here is how to assign values of the fitted parameters and other relevant input parameters (e.g., IR photon temperature and luminosity, etc.) to each member of the UHECR-neutrino emitting TDE population. We are taking a model-independent approach and assuming that each of the AT2019aalc and AT2019fdr represents its own unique group. We will refer to them as the aalc-group and the fdr-group, which together represent the total population. Each group has its unique parameters and local rate (, where aalc or fdr). The local rate of the entire UHECR-neutrino emitting TDE population is given by . The analyzed range and values of these parameters are summarized in Tab. 4
@@
Considering the large number of fitting variables and the limited amount of data, we have to consider different scenarios, which are compared to each other. In particular, we examine the following:
@@
For each scenario, we are considering two options for the composition of the injected CR, as discussed in Sec. 2.3. The first option follows the maximum abundance of a star from the selected statistics. The second option treats the composition of the injected CR as fitting parameters.
@@ 3.3 UHECR transport to Earth and description of UHECR data
In our analysis of UHECRs originating from TDEs, we employ the open-source software PriNCe (J. Heinze et al., 2019), designed to numerically solve the one-dimensional transport equations, to model the propagation of CRs in the Universe and compute the cumulative cosmogenic neutrino spectra. We propagate the CRs escaped from the source for each aalc- and fdr-group, considering varying radius, maximal rigidity, and injected composition. We consider a redshift evolution up to for the TDE populations to calculate the cumulative diffuse neutrino fluxes. For the configuration of the calculation, we use the build-in Talys (A. J. Koning et al., 2007) nuclear interaction model, and use the Extragalactic Background Light (EBL) model by R. C. Gilmore et al. (2012).
Our primary goal is to fit the UHECR spectrum (V. Verzi, 2019) and composition (A. Yushkov, 2019) data above , observed by the Pierre Auger Observatory (Auger). In addition, we ensure that the diffuse neutrino flux from the TDE population does not exceed the High-Energy Starting Events (HESE) data from IceCube. For the composition fitting, we attempt to reproduce the observed mean depth of shower maximum density () and its standard deviation (), using the EPOS-LHC air-shower model (T. Pierog et al., 2015)444We also tested the Sibyll 2.3d air-shower model (F. Riehn et al., 2020) and found no significant difference. The results from Sibyll 2.3d air-show model are available at plotkopavlo.com/uhecrsources/tde.. To account for systematic, we include the energy and uncertainties. However, we exclude the uncertainties since they have a minor influence on our analysis.
To evaluate the goodness of the fitting, we adopt the method from P. Plotko et al. (2023) and calculate the value for the corresponding model of the TDE population:
@@
Further, to compare different models of the TDE population, we compute the Akaike Information Criterion corrected (, H. Akaike, 1974; K. P. Burnham & D. R. Anderson, 2004; Buchner, J. et al., 2014; A. Rosales et al., 2020) through the given equation:
where is the number of data points, and is the number of parameters of the corresponding model . The model with the smallest value explains the data best among tested models. It is important to highlight that should not be used as a null-hypothesis test (D. R. Anderson et al., 2000; R. Mundry, 2011). In our study, we compare the score of each model against of the model with the smallest value (), calculating the difference () as:
To make it easier to observe the differences between models, we compute in the units of standard deviations following P. Plotko et al. (2023), and use it as an identifier for model selections.
@@ 4.1 Evaluation of TDE population models
. Scenario Composition Group’s local rate Fit quality metrics aalc-group fdr-group aalc-group fdr-group / d.o.f. Mono aalc MS 10 654/29 603 CO-WD 59 84/29 33 5.3 ONeMg-WD 75 192/29 140 Fitted 122 57/27 12 2.8 fdr Fitted 9 261/27 216 Democratic MS 69 37650/29 37599 CO-WD 75 2420/29 2369 ONeMg-WD 69 447/29 396 Fitted 69 248/27 203 Diverse ONeMg-WD WR 47 5 57/24 23 4.3 Fitted Fitted 194 13 38/25 Best-fit case
@@ 4.2 Describing the UHECR spectra and composition with TDE population
In this section, we present results for the best-fit case of the UHECR-neutrino emitting TDE population, which is the Diverse Scenario in the Fitted Composition Case. In this scenario, the aalc- and fdr-groups have different parameters, such as the radius of the production region, the maximum rigidity, and the injected CR composition. Fig. 6 illustrates the predictions for UHECR spectrum and composition and diffuse neutrino flux with a comparison to the data for the best-fit model. The obtained fitting parameters for the Diverse Scenario are detailed in Tab. 6.
@@
Our analysis shows that the predicted UHECR spectrum closely aligns with the spectrum observed by the Auger Observatory, with the fitted value of the systematic energy shift of about -9.4%. The aalc-group primarily dominates the UHECR spectrum. Although the fdr-group’s overall contribution is subdominant in spectral fitting, it plays a crucial role in composition fitting, particularly at the highest energies. The model adjustment results in a heavier composition by shifting by approximately 174%. Notably, this adjustment is consistent with recent findings from Auger, e.g., a potential systematic bias toward a heavier composition (A. Abdul Halim et al., 2024b).
@@
In this scenario, the diffuse source neutrino flux from the TDE population does not exceed IceCube’s HESE data. The contribution becomes significant at energies of GeV and higher, peaking at GeV. Unlike the UHECR spectrum, both the aalc- and fdr- groups contribute at a comparable level to the diffuse neutrino background (see the upper right panel of Fig. 6). However, we also observe that the aalc-group produces a more peaked spectrum with a lower maximum energy in contrast to the fdr-group. As a result, the two groups predict different slopes of the neutrino spectrum in the EeV energy range. As can be seen on the upper right panel of Fig. 6), both groups are likely detectable by forthcoming neutrino detectors like IceCube-Gen2 and GRAND200k. Regrettably, the predicted diffuse flux is too low for RNO-G to detect.
@@
Regarding the composition of injected UHECR, the aalc-group emits CRs in both intermediate-mass (e.g., sodium or silicon) and the light (e.g., hydrogen and helium) ranges, whereas the fdr-group mostly contributes the heavy nuclei in the best-fit scenario. This distribution is depicted in Fig. 7 and detailed in Tab. 6. The composition fitting for the aalc-group is similar to ONeMg-WD. Still, it requires additional light elements. The absence of intermediate elements could be attributed to the loss of intermediate elements before they are transferred to the acceleration zone. Furthermore, more irons are needed for the fdr-group to make the predicted composition heavier and better fit the composition data. However, within the 1 AICc region, the injected composition could be pure silicon, which is similar to the ONeMg-WD. One way to achieve an iron-abundant CR injection for both groups is by disrupting a MS star with an enhancement factor for the acceleration of heavier elements, which will be discussed in the Sec. 5.
@@
In order to match the observed flux of UHECR, a local rate of 200 is required for the aalc-group. This value is typical for TDEs that disrupt MS stars rather than white dwarfs. For the fdr-group, the local rate should be around 10 , which is more indicative for post-main-sequence stars.
Further, Fig. 8 explores the parameter spaces for the aalc- and fdr-groups, and Fig. 9 presents the escaped spectra of CRs and neutrinos for the best-fit parameters. The aalc-group parameters are notably similar to those previously modeled for AT2019aalc (W. Winter & C. Lunardini, 2023), with a maximum rigidity value of GV that is expected for UHECR sources (J. Heinze et al., 2019; A. Aab et al., 2017; P. Plotko et al., 2023). The composition 3 region requires values between and GV to fit the data. Higher values are excluded since, in this case, the escape CRs from TDE are not dominated by a single element group but also have too many secondary light elements, contradicting the data.
On the other hand, the 3 allowed region, based on the spectral fitting, for the aalc-group (marked as the yellow region in the left panel of Fig. 8) is broad and does not significantly impact the fit. The spectrum is less sensitive to composition and does not provide a clear reason for this shape of the 3 spectrum allowed region. The required value of the TDE source radius of the aalc-group is cm, falling within the purple region allowed by the neutrino constraints for AT2019aalc.
Conversely, the fdr-group requires substantially different parameters from previously used in W. Winter & C. Lunardini (2023). In order to fit the composition data, the fdr-group should have a higher maximum rigidity, with the best value of GV. This suggests that the TDEs from the fdr-group could be a powerful UHECR source capable of accelerating iron to energies of GeV. The preference for a high maximum rigidity arises because iron nuclei are required to contribute at energies around GeV, while their presence at lower energies should be negligible. Spectral constraints are not applied to the fdr-group since its overall contribution to the UHECR spectrum is too low to significantly impact the fit.
One may wonder why in Fig. 6 the UHECR flux for the aalc-group is significantly higher than that of the fdr-group (upper left panel), whereas the neutrino fluxes are comparable (upper right panel). As one can see in Fig. 9, fdr-like TDEs produce a (relatively) higher neutrino flux (compare dashed/dashed-dotted curves with solid black curves). The reason is the higher optical thickness due to interactions, driven by the higher luminosities (see Tab. 1) in combination with the much smaller (see Tab. 6).
@@ 5 Enhancement of heavier nuclei in the acceleration process
In order to further investigate the possibility of obtaining a heavier CR composition, we introduce an enhancement factor for the acceleration of heavier elements. This option is attractive in combination with the very abundant MS disruptions, which however exhibit a light progentitor composition. Realistically, nucleus selection effects may exist and cause the composition of the accelerated nuclei to differ from that of the disrupted star. One such effect might be partial disruption, which is supported by recent observations (J. J. Somalwar et al., 2023). In this case, the core of the star would remain intact while the matter in the outer layers is disrupted and subsequently accreted onto the central black hole. Ultimately, the accreted material could be accelerated, resulting in the composition of the observed UHECRs. Another possibility is that there might be enhancement mechanisms in the acceleration process that would favor heavier nuclei, thus skewing the UHECR toward heavier composition compared to the original star.
We define an enhancement factor based on D. Caprioli et al. (2017) and A. Hanusch et al. (2019), as D. Caprioli et al. (2017) showed that nonrelativistic collisionless shocks can preferentially accelerate ions based on their mass/charge ratio ( ratios). As a result, these shocks develop a non-thermal tail whose normalization is enhanced as . A. Hanusch et al. (2019) further supports the acceleration of heavier ions through hybrid simulations, which reveal that the injection efficiency of ions into the diffusive shock acceleration process depends strongly on the shock Mach number. Their simulations showed that the injection efficiency for ions with high ratios is enhanced. To generalize the idea, we assume that injection efficiency is proportional to , where is the enhancement factor. Additionally, we assume that thermal particles lose an electron () before being transferred to an acceleration zone to maximize the effect. The net effect is that the abundance of the element in the accelerated stream is scaled by a power of the mass number, leading to new, rescaled mass fractions, :
@@
In our analysis, we use the Diverse Scenario and focus on the fixed composition of a MS star with various values of for injected CRs. In this scenario, we assume that the aalc- and fdr-groups have different values. For each combination of and , we conduct a fit to determine the minimum value while minimizing over all other fitted parameters. We show as a function of the enhancement factor in Fig. 11. For the aalc-group, the fit requires the value of to be between 1.5 and 2. For the fdr-group, there is a slight preference for values of for the fdr-group, but not statistically significant due to the small contribution of the fdr-group to the UHECR spectrum.
Overall, the Enhancement Factor Case is excluded by more than 8 compared to the Fitted Composition Case due to the nature of the enhancement factor and the assumed fixed composition. The model predicts a light composition for values of the enhancement factor near zero, which contradicts the UHECR composition observations. As the value of increases, the model becomes too efficient at accelerating heavy elements. However, to explain UHECR data, an intermediate-mass group is needed. Different assumptions could help address this issue. For example, there may be two groups of TDEs: one that disrupts white dwarfs with intermediate mass elements and another that disrupts MS stars with an enhancement factor for acceleration of heavier elements. This approach would provide both lighter and heavier elements, potentially improving the fit of the composition data. However, such a model would introduce too many free parameters and could be statistically excluded in favor of the simpler Fitted Composition Case.
Fig. 12 compares the best-fit case to the Auger spectrum and composition, along with the predicted diffuse neutrino flux, for the Enhancement Factor Case. The corresponding parameters of the best-fit case are presented in Tab. 7. While the model’s predictions can explain the UHECR spectrum, the main contribution comes from the iron group, which contradicts the current UHECR composition observation. Additionally, the diffuse neutrino flux from the aalc-group exceeds the HESE limits to a high local rate of 1116 . These two factors result in a poor quality of fit. One potential solution to improve the fit could be using a more complex enhancement factor, such as having iron elements lose more electrons than the silicon group, resulting in a more silicon-dominant composition.